slope intercept form
a. m= -6; (9,3)
b. m= -2 b=4

Answer by Alan3354(69443)

slope intercept form
a. m= -6; (9,3)
--------------
y-y1 = m*(x-x1)
y-3 = -6*(x-9)
y-3 = -6x + 54
y = -6x + 57 (slope-intercept form)
6x+y = -57 (standard form)
b. m= -2 b=4
y = mx + b
y = -2x + 4 SI form
2x+y = 4 standard form